OpenStudy (anonymous):

Determine the dimensions of a rectangular solid (with a square base) with maximum volume if its surface area is 100 meters.

OpenStudy (ranga):

Let x be the side of the square base and h be the height. Volume V = ? Surface Area = ?

OpenStudy (anonymous):

well the surface area is 100

OpenStudy (ranga):

Yeah, but what is the surface area in terms of x and h? what is the volume in terms of x and h?

OpenStudy (ranga):

A rectangular solid with a square base has 2 squares (top square & bottom square). It has 4 rectangles (4 sides) Add up the areas of 2 squares of side x and 4 rectangles of sides x and h.

OpenStudy (anonymous):

v=x(h) can't think of sa off the top of my head right so im going to guess its sa=0.5(x)(h)?

OpenStudy (ranga):

A rectangular solid is a prism. Volume of a prism = area of base * height

OpenStudy (ranga):



OpenStudy (ranga):

Volume V = x^2 * h ----- (1) Surface area = 2 * x^2 + 4 * x * h = 100 ------ (2) From (2), find h. Substitute h in (1). Now you will have V as a function of just one variable, namely, x. To maximize V, find dV/dx, equate it to zero and solve for x.

OpenStudy (anonymous):

ok ill work on it. thank you!

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

ok @ranga i got base side=(10√6)/3 and height=(5√6)/3

OpenStudy (ranga):

I am getting x = (5√6)/3. But I will have to cjhck my calculations again.

OpenStudy (anonymous):

i was getting all kinds of things for x and (5√6)/3. just rechecked it again and thats what i got. did we get the same thing for h though?

OpenStudy (ranga):

I am getting x = (5√6)/3 and h = (5√6)/3. So the solid turns out to be a cube which seems right to me.

OpenStudy (anonymous):

ok